OpenStudy (anonymous):

i Need Help On Geometry, WHat aree THee Intercepts for the line 3x + 4y = 7

OpenStudy (anonymous):

1step let x=0 find y (0,y)- it is 1st intercept 2step let y=0 find x (x,0)- it is 2nd intercept

OpenStudy (anonymous):

Ooo, Okaii One more Can u Help Me, what is the equation of a line that passes through (5, 3) and (-5, 3)

OpenStudy (anonymous):

try y-y1=m(x-x1) got idea?

OpenStudy (anonymous):

But That WOuld ii Put as y and y1?

OpenStudy (anonymous):

didn't pay attention. This case even easier. It's horizontal line. try to plot both points

OpenStudy (anonymous):

Haha Kay :) THanksz

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i dont get it its hard :/ ii tryed both bt the answers are y=3, y= -3., x = 5 and x= -5

OpenStudy (anonymous):

for the 2nd ? answer is y=3

OpenStudy (anonymous):

But how u get that thoe?

OpenStudy (anonymous):